OpenStudy (fantino_italiano):

New, Please help me!!! What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)? A. y − 5 = 3(x − 1) B. y − 2 = −3(x − 0) C. y + 2 = 3(x + 1) D. y + 5 = −3(x + 1)

OpenStudy (fantino_italiano):

Someone help lol

OpenStudy (fantino_italiano):

okay man like I need help, and I just pressed a "bump" button I don't know what that means lol

OpenStudy (kl0723):

\[y – y_{1} = m(x – x_{2})\]

OpenStudy (kl0723):

this is your slope form equation, means the answer should be in this form

OpenStudy (kl0723):

now we need to identify which one is which and find the slope which is "m"

OpenStudy (kl0723):

first point is (0,2) second point is (1,5)

OpenStudy (fantino_italiano):

okay awesome

OpenStudy (kl0723):

x1=0 y1=2

OpenStudy (kl0723):

that's for the first point

OpenStudy (fantino_italiano):

Okay

OpenStudy (kl0723):

then we have that x2=1 and y2=5

OpenStudy (kl0723):

now you can use the slope formula: \[m=\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (kl0723):

can you plug in the numbers and solve for it?

OpenStudy (kl0723):

what would be the result?

OpenStudy (fantino_italiano):

Hey sorry lol this whole site is confusing, but is the answer d? That was my initial thought

OpenStudy (kl0723):

m=3

OpenStudy (kl0723):

plug in all the values into the equation form to get your answer

OpenStudy (fantino_italiano):

Okay so like M=y(5-1) And then x (0- 2)?

OpenStudy (fantino_italiano):

Wait no that's wrong I wrote it wrong one sec

OpenStudy (kl0723):

this is the equation: y–y1=m(x–x2) you have y1=2 and x2=1

OpenStudy (kl0723):

now we plug in the values tht we know, the rest is the same: y-2 = 3(x-1)

OpenStudy (fantino_italiano):

Can you just message me this site keeps redirecting me

OpenStudy (kl0723):

I know, it's kind of buggy

OpenStudy (fantino_italiano):

Yeah message me real quick

OpenStudy (kl0723):

i did